# Discrete Cosine Transform Based Decorrelated Attention for Vision Transformers

### v3

@@ III-B DCT-based Initialization for Attention
III-C DCT-Based Compressed Attention
Data compression using DCT is common in signal, image, and video processing for energy concentration, compression efficiency, and noise reduction. DCT not only decorrelates the input image data similar to the Karhunen-Loeve Transform (KLT) or principal component analysis (PCA) but also has a frequency interpretation because it can be computed using the discrete Fourier transform (DFT) [41]:
$\begin{split}&\{\mathbf{X}\mathcal{D}^{T}\}[k]\\ &=\begin{cases}\frac{\sqrt{2}}{C}\sum_{n=0}^{C-1}\mathbf{X}[n],&k=0,\\ \frac{2}{C}\text{Re}\left(e^{-\frac{ik\pi}{2C}}\sum_{n=0}^{2C-1}\mathbf{X}[n]W^{kn}\right),&k=1,\ldots,C-1,\end{cases}\end{split}$ (5)
@@ III-C DCT-Based Compressed Attention
where $W=e^{-\frac{i2\pi}{2C}}$, $i^{2}=-1$, and $\text{Re}(\cdot)$ implies the real part of the term enclosed. Usually, higher-frequency DCT coefficients contribute less to perceptual quality and contain more noise. Therefore, preserving essential information in lower frequencies while discarding less critical details in higher frequencies results in efficient signal representation and storage. The specific truncation strategy depends on the application’s requirements and the acceptable level of information loss.
@@
We observed that the attention layer matrices are correlated with each other so that we can compress them using DCT. Fig. 1b shows the method of DCT-based compression for the attention layer. First, the input patches $\mathbf{X}\in\mathbb{R}^{N\times M^{2}\times C}$ are encoded by DCT along the channel of $C$ with the high-frequency components truncated. If we keep the first $\tau C$ coefficients ($\tau<1$), the essential information of $\mathbf{X}$ is still preserved, while the input and output dimensions of the linear layers ($\mathcal{L}_{Q},\mathcal{L}_{K},$ and $\mathcal{L}_{V}$) are reduced from $\mathbb{R}^{N\times M^{2}\times C}$ to $\mathbb{R}^{N\times M^{2}\times\tau C}$. In this way, the computational cost from $\mathcal{L}_{Q},\mathcal{L}_{K}$, and $\mathcal{L}_{V}$ is reduced significantly. In our experiments, $\tau$ is chosen from 25%, 50%, and 75%. After obtaining the attention $\mathbf{A}\in\mathbb{R}^{N\times M^{2}\times\tau C}$ from the queries $\mathbf{Q}$, keys $\mathbf{K}$, and values $\mathbf{V}$, we apply zero-padding along the last channel to make the dimension back to $\mathbb{R}^{N\times M^{2}\times C}$. Before being fed into the linear layer $\mathcal{L}_{O}$, these padded patches are decoded by the inverse DCT (IDCT):
$\begin{split}\{\mathbf{A}(\mathcal{D}^{-1})^{T}\}[n]=\frac{\sqrt{2}}{2}\mathbf{A}[0]+\sum_{k=1}^{C-1}\mathbf{A}[k]\cos\frac{(2n+1)k\pi}{2C}&,\\ n=0,1,\ldots,C-1&.\end{split}$ (6)
@@
Because $\mathcal{D}$ is orthogonal, i.e. $\mathcal{DD}^{T}=\mathbf{I}_{C\times C}$, the IDCT matrix can be obtained directly from the transpose of the DCT matrix. DCT and IDCT can be implemented in a fast manner with the complexity of $O(C\log_{2}C)$ using butterfly algorithm [42]. However, such an implementation is not officially supported in PyTorch currently. Therefore, in this work, we implement the DCT and IDCT via matrix multiplication between the input tensor and the truncated DCT and IDCT matrices $\bar{\mathcal{D}}\in\mathbb{R}^{\tau C\times C}$ and $\bar{\mathcal{D}}^{-1}\in\mathbb{R}^{C\times\tau C}$. These transform matrices are obtained by truncating the DCT and IDCT matrices along rows or columns to keep $\tau$ coefficients or zero padding, respectively. Then, DCT with truncation or IDCT with zero padding is implemented as one matrix multiplication to reduce memory pressure. Furthermore, as no non-linearity exists between IDCT and the final linear layer $\mathcal{L}_{O}$, these two steps can be implemented as a single linear layer whose weight matrix is $\mathbf{W}_{O}\bar{\mathcal{D}}^{-1}$ and the bias is still $\mathbf{b}_{O}$. However, such a combination is not applied on DCT and $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, or $\mathbf{W}_{V}$. This is because it may not always reduce the computational cost for them: To compute $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$, with such a combination, $3kNM^{2}C^{3}$ multiplications are needed. In contrast, if they are computed separately, there are $\tau NM^{2}C^{3}$ multiplications from DCT and $\tau^{2}NM^{2}C^{3}$ multiplications from each linear layer. As a result, there are a total of $(\tau+3\tau^{2})NM^{2}C^{3}$ multiplications. Hence, combing DCT with $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$ and $\mathbf{W}_{V}$ brings computational cost saving only when $\tau>\frac{2}{3}$. To maintain the consistency of the framework, we do not combine DCT with $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, or $\mathbf{W}_{V}$.
Computational cost comparison of the vanilla versus the proposed DCT-compressed attentions is presented in Table I. We evaluate both naive and simplified implementations of the DCT-compressed attention with the vanilla attention. The simplification is implemented by fusing IDCT into the linear layer $\mathcal{L}_{O}$. Computational overhead from the Softmax function is omitted because the Softmax function is based on exponential terms. Despite this, the Softmax function in the DCT-compressed attention requires less computational cost than in the vanilla attention as the number of entries in $\left(\frac{\mathbf{QK}^{T}}{\sqrt{d}}+\mathbf{B}\right)$ in the DCT-compressed attention is $\tau^{2}$ times ($\tau<1$) as in the vanilla attention.
TABLE I: Multiplications in MSA on $\mathbf{X}\in\mathbb{R}^{N\times M^{2}\times C}$. “--” means this step is not required. $\tau<1$.
@@
$\mathbf{QK}^{T}/\sqrt{d}+\mathbf{B}$ $NM^{4}C^{2}$ $\tau^{2}NM^{4}C^{2}$ $\tau^{2}NM^{4}C^{2}$
Multiply with $\mathbf{V}$ $NM^{4}C^{2}$ $\tau^{2}NM^{4}C^{2}$ $\tau^{2}NM^{4}C^{2}$
IDCT - $\tau NM^{2}C^{3}$
@@
Total $4NM^{2}C^{3}+2NM^{4}C^{2}$ $(1+2\tau+3\tau^{2})NM^{2}C^{3}+2\tau^{2}NM^{4}C^{2}$ $(2\tau+3\tau^{2})NM^{2}C^{3}+2\tau^{2}NM^{4}C^{2}$
III-D Frequency Domain Motivation of the DCT-based Initialization
In traditional random initialization, all the weight matrices are initialized as random numbers that are independently generated. As a result, the initial matrices can be considered as a realization of white noise covering the entire spectrum. A good initialization should take advantage of the full bandwidth to cover the entire spectrum of the input.
The DCT-based initialization also covers the entire frequency domain. The set of DCT basis vectors $\{\frac{\sqrt{2}}{2},\cos\frac{(2n+1)k\pi}{2C}\}$ is the class of discrete Chebyshev polynomials:
@@ III-D Frequency Domain Motivation of the DCT-based Initialization
for $n=0,1,\ldots,C-1$. The basis member $\cos\frac{(2n+1)k\pi}{2nC}$ is the $k$-th Chebyshev polynomial $T_{k}[\xi]$ evaluated at the $n$-th zero of $T_{n}[\xi]$ [41]. Fig. 2 draws the basis vectors in Eq. (7) for an $8\times 8$ DCT matrix and Fig. 3 presents the magnitude of their discrete Fourier transform (${}\mathcal{F}\{\cdot\}$):
@@
Figure 3: Frequency response of the DCT basis vectors. Each frequency response is shown in a solid curve in different colors. The frequency is normalized to $[0,\pi]$. The black dashed line denotes the root of the squared sum of these frequency responses, showing the entire spectrum is covered.
III-E Motivation of the DCT-based Compression
The mechanism behind the DCT-based compression mimics the Karhunen-Loeve transform (KLT), which is also known as principal component analysis (PCA), a mathematical technique used for dimensionality reduction in signal processing [43, 44]. Neighboring image patches are highly correlated with each other. During training, the attention matrices are inter-correlated because they are obtained using similar gradient vectors. Therefore, they can be decorrelated and compressed using DCT. When the input data is highly correlated, the covariance matrix can be approximated by a Teoplitz matrix $\phi$ of $[1,\rho,\rho^{2},\cdots,\rho^{C-1}]$:
@@ III-E Motivation of the DCT-based Compression
with $\rho<1$ close to 1 [45]. This is the covariance matrix of an autoregressive model random process AR(1) with the correlation coefficient $\rho$ [46, 47]. Fig. 2 shows that DCT basis vectors approximate the KLT of a first-order Markov process with high correlation among the neighboring samples [48, 49]. We observe that $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ have parameters with high correlations among themselves. We can directly use DCT in data compression since its matrix approximates the corresponding KLT matrix [41],
Although KLT can also be utilized in data compression, it requires massive computation for the covariance matrices and the corresponding eigenvectors of $\mathbf{X}$. On the other hand, DCT is very successful in decorrelating the data without computing the covariance matrix and eigenvectors. As a consequence, it is also used in JPEG image [50, 51] and MPEG video [52] compression methods.
IV Experimental Results
@@
We conduct experiments on the CIFAR-10 [53] and ImageNet-1K [54] classification tasks and COCO 2017 detection and segmentation tasks [55]. Swin Transformers [17] are employed as the backbone network. Models are trained on a server with 8 NVIDIA A100 80GB GPUs using Python with PyTorch framework [56].
IV-A Implementation Details
IV-A1 Image Classification Tasks Implementation Details
@@
Training models for the image classification tasks from scratch mostly follow the settings in [17]. We employ AdamW optimizer [57] for 300 epochs using a cosine decay learning rate scheduler and 20 epochs of linear warm-up. We set the batch size as 1024, the initial learning rate as 0.001, and the weight decay as 0.05. Data augmentation is performed as the following: The CIFAR-10 augmentation follows [58]. First, the training images are padded with 4 pixels. Then, they are randomly cropped to get 32 by 32. Finally, the images are randomly flipped horizontally. The images are normalized with the means of [0.4914, 0.4822, 0.4465] and the standard variations of [0.2023, 0.1994, 02010]. The ImageNet-1K augmentation follows PyTorch Torchvision’s GitHub repository [56] [^1] https://github.com/pytorch/vision/tree/release/2.0/references/classification. As a start, images are resized to make the smaller edge of the image 232 for Swin-T and 246 for Swin-S. Next, they are cropped to 224 by 224 images and randomly flipped horizontally. Techniques such as Autoaugment [59], mixup [60], cutmix [61], and random erasing [62] are applied. Images are normalized with the means of [0.485, 0.456, 0.406] and the standard deviations of [0.229, 0.224, 0.225]. During the training, the best models are stored based on the top-1 accuracy on the CIFAR-10 test dataset and the ImageNet-1K validation dataset.
Flops computation for the Swin Transformers and our revisions uses the function “flops” in the official implementation of [17] [^2] https://github.com/microsoft/Swin-Transformer/blob/main/models/swin_transformer.py. To follow their work, the flops from the Softmax function are omitted. Flops computation for the ViT-B-32 and our revisions uses Python library “ptflops” [^3] https://github.com/sovrasov/flops-counter.pytorch, with manually adding the flops from DCT.
IV-A2 COCO Detection and Segmentation Tasks Implementation Details
Training models for the COCO tasks follows the settings in [17]. Models start from ImageNet-1K pre-trained weights. We use AdamW optimizer [57] for 36 epochs with an initial learning rate of 0.0001, weight decay of 0.05, batch size of 16, 3$\times$ schedule, and multi-scale training [63, 64].
IV-B CIFAR-10 Classification
Training vision transformers from scratch on the CIFAR-10 dataset is notoriously difficult. We demonstrate the benefits of using our DCT initialization and DCT compression for the Swin Transformer Tiny (Swin-T) on from-scratch CIFAR-10 training to show the effectiveness of our methods on the small-scale dataset. Table II presents the performance comparison on the CIFAR-10 test dataset, where the number of parameters only counts trainable parameters.
@@ IV-B CIFAR-10 Classification
TABLE II: CIFAR-10 test dataset evaluation. In column “DCT”, Q/K/V denotes the weights initialized as the DCT matrix, and numbers mean the proportion of coefficients kept in DCT-compressed attention. Models with higher or lower accuracy are highlighted in pink or cyan, respectively.
Model DCT Params (M) Top-1 Acc.
Swin-T - 27.53 90.51%
Swin-T Q 27.53 90.78%
Swin-T K 27.53 90.83%
Swin-T V 27.53 90.81%
Swin-T 0.25 21.45 90.46%
Swin-T 0.5 22.67 90.67%
Swin-T 0.75 24.69 90.90%
We start by training Swin-T with the vanilla attention function for the baseline and obtain 90.51% accuracy on the CIFAR-10 test dataset. Then, we investigate the DCT-initialized attention in Swin-T on the CIFAR datasets. The accuracy of Swin-T on the CIFAR-10 dataset is always improved regardless of which attention weight is initialized as the DCT matrix. When the weights to compute keys ($\mathbf{K}$) are initialized as the DCT matrices, the accuracy is improved to 90.83%.
@@
Next, we evaluate the DCT-compressed attention with different truncating rates in Swin-T on the CIFAR-10 dataset. Table II presents when we only keep 25% of DCT coefficients, the number of trainable parameters is reduced from 27.53M to 21.45M (22.1% off), while the accuracy only drops 0.05%. If we keep more DCT coefficients, the accuracy even improves. If we keep 75% of DCT coefficients, the accuracy increases from 90.51% to 90.90%. This is because although the number of trainable parameters is reduced, transforming the input data into the frequency domain using DCT can boost the extraction of important features. As a result, fewer trainable parameters are required.
IV-C ImageNet-1K Classification
ImageNet-1K [54] is a large-scale visual database designed for visual object recognition research. It was created to facilitate the development, training, and testing of computer vision models, especially in deep learning. ImageNet consists of a vast collection of labeled images, covering a wide range of object categories.
Table III presents the performance comparison on the ImageNet-1K dataset, where the number of parameters only counts trainable parameters. First, we investigate the DCT-initialized attention in the Swin-T and the Swin-S on the ImageNet-1K dataset. The accuracy on the ImageNet-1K validation dataset is always improved nevertheless which weight matrix is initialized as the DCT matrix. For example, when we initialize the weights to compute keys ($\mathbf{K}$) as the DCT weights, the accuracy of Swin-T from 81.474% to 81.728% and Swin-S are improved from 83.196% to 83.378%, respectively.
Then, we evaluate the DCT-compressed attention with different truncating rates and compare the results with DCFormer [36], as it is also a DCT-based method to reduce the computational cost. Table III presents that when we keep 75% of DCT coefficients, we can obtain a comparable accuracy result on Swin-T and a slightly better accuracy result on Swin-S. Contrarily, the trainable parameters of the Swin-T and the Swin-S are reduced from 28.29M to 25.45M (13.1% off) and from 49.61M to 44.45M (10.4% off), respectively. Flops are reduced from 4.49G to 4.18G (6.9% off) and 8.13G to 8.74G (7.0% off), respectively. If fewer DCT coefficients are kept, the swim transformers can get further savings on parameters and computational overhead. Meanwhile, the accuracy will drop correspondingly. As a comparison, although DCFormer-SW-T/S [36] with $\tau=2$ reduce the FLOPs significantly ($\tau=2$ means 4×4$\times$4 × frequency compression), the performance drops seriously because the DCFormer takes both low-pass and high-pass DCT components equivalent important.
We further evaluate the two proposed DCT-based attentions in the ViT-B-32 backbone [16] to verify their performance on other vision transformers. When we initialize the weights to compute keys ($\mathbf{K}$) as the DCT weights, the accuracy of ViT-B-32 is improved from 73.120% to 73.614%. When we apply DCT-compressed attention with 75% coefficients kept, the accuracy increases from 73.120% to 73.226% while the number of trainable parameters and FLOPs decline from 88.22M to 75.83M (14.0% off) and reduced from 4.41G to 4.05G (8.2% off), respectively.
TABLE III: ImageNet-1K validation dataset evaluation. After DCT, Q/K/V denotes the weights initialized as the DCT matrix, and numbers mean the proportion of coefficients kept in DCT-compressed attention. Our models with higher or lower top-1 accuracy are highlighted in pink or cyan, respectively.
Method Params (M) Flops (G) Top-1 Acc. Top-5 Acc.
Swin-T (Torchvision) [17] 28.29 4.49 81.474% 95.776%
PoolFormer-S24 [40] 21 3.5 80.3% -
DCFormer-SW-T ($\tau=1)$ [36] 28.29 4.5 81.2% -
DCFormer-SW-T ($\tau=2)$ [36] 28.29 1.3 79.2% -
gSwin-T [20] 22 3.6 81.715% -
@@ IV-C ImageNet-1K Classification
Swin-S (Torchvision) [17] 49.61 8.74 83.196% 96.360%
PoolFormer-M36 [40] 56 8.8 82.1% -
DCFormer-SW-S ($\tau=1)$ [36] 49.61 8.7 82.8% -
DCFormer-SW-S ($\tau=2)$ [36] 49.61 2.7 80.9% -
gSwin-S [20] 40 7.0 83.014% -

### v5

@@ 3.2 DCT-based Initialization for Attention
3.3 DCT-Based Compressed Attention
@@
Data compression using DCT is common in signal, image, and video processing for energy concentration, compression efficiency, and noise reduction. Usually, higher-frequency DCT coefficients contribute less to perceptual quality and contain more noise. Therefore, preserving essential information in lower frequencies while discarding less critical details in higher frequencies results in efficient signal representation and storage. The specific truncation strategy depends on the application’s requirements and the acceptable level of information loss.
@@
We observed that the attention layer matrices are correlated with each other, so that we can compress them using DCT. Fig. 1b shows the method of DCT-based compression for the attention layer. First, the input patches $\mathbf{X}\in\mathbb{R}^{N\times M^{2}\times C}$ are encoded by DCT along the channel of $C$ with the high-frequency components truncated. If we keep the first $\tau C$ coefficients ($\tau<1$), the essential information of $\mathbf{X}$ is still preserved, while the input and output dimensions of the linear layers ($\mathcal{L}_{Q},\mathcal{L}_{K},$ and $\mathcal{L}_{V}$) are reduced from $\mathbb{R}^{N\times M^{2}\times C}$ to $\mathbb{R}^{N\times M^{2}\times\tau C}$. In this way, the computational cost from $\mathcal{L}_{Q},\mathcal{L}_{K}$, and $\mathcal{L}_{V}$ is reduced significantly. In our experiments, $\tau$ is chosen from 25%, 50%, and 75%. After obtaining the attention $\mathbf{A}\in\mathbb{R}^{N\times M^{2}\times\tau C}$ from the queries $\mathbf{Q}$, keys $\mathbf{K}$, and values $\mathbf{V}$, we apply zero-padding along the last channel to make the dimension back to $\mathbb{R}^{N\times M^{2}\times C}$. Before being fed into the linear layer $\mathcal{L}_{O}$, these padded patches are decoded by the inverse DCT (IDCT):
[…]
@@
Because $\mathcal{D}$ is orthogonal, i.e. $\mathcal{DD}^{T}=\mathbf{I}_{C\times C}$, the IDCT matrix can be obtained directly from the transpose of the DCT matrix. Although DCT and IDCT can be implemented in a fast manner with the complexity of $O(C\log_{2}C)$ using the butterfly algorithm Chen et al. (1977), such an implementation is not officially supported in PyTorch currently. Therefore, in this work, we implement the DCT and IDCT via matrix multiplication between the input tensor and the truncated DCT and IDCT matrices $\bar{\mathcal{D}}\in\mathbb{R}^{\tau C\times C}$ and $\bar{\mathcal{D}}^{-1}\in\mathbb{R}^{C\times\tau C}$. These transform matrices are obtained by truncating the DCT and IDCT matrices along rows or columns to keep $\tau$ coefficients or zero padding, respectively. Then, DCT with truncation or IDCT with zero padding is implemented as one matrix multiplication to reduce memory pressure. Furthermore, as no non-linearity exists between IDCT and the final linear layer $\mathcal{L}_{O}$, these two steps can be implemented as a single linear layer whose weight matrix is $\mathbf{W}_{O}\bar{\mathcal{D}}^{-1}$ and the bias is still $\mathbf{b}_{O}$. However, such a combination is not applied on DCT and $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, or $\mathbf{W}_{V}$. This is because it may not always reduce the computational cost for them: To compute $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$, with such a combination, $3kNM^{2}C^{3}$ multiplications are needed. In contrast, if they are computed separately, there are $\tau NM^{2}C^{3}$ multiplications from DCT and $\tau^{2}NM^{2}C^{3}$ multiplications from each linear layer. As a result, there are a total of $(\tau+3\tau^{2})NM^{2}C^{3}$ multiplications. Hence, combing DCT with $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$ and $\mathbf{W}_{V}$ brings computational cost saving only when $\tau>\frac{2}{3}$. To maintain the consistency of the framework, we do not combine DCT with $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, or $\mathbf{W}_{V}$.
@@
$\frac{\mathbf{QK}^{T}}{\sqrt{d}}+\mathbf{B}$ $NM^{4}C^{2}$ $\tau^{2}NM^{4}C^{2}$ $\tau^{2}NM^{4}C^{2}$
Multiply $\mathbf{V}$ $NM^{4}C^{2}$ $\tau^{2}NM^{4}C^{2}$ $\tau^{2}NM^{4}C^{2}$
IDCT - $\tau NM^{2}C^{3}$ -
@@
Total $4NM^{2}C^{3}$ $(1+2\tau+3\tau^{2})NM^{2}C^{3}$ $(2\tau+3\tau^{2})NM^{2}C^{3}$
$2NM^{4}C^{2} +$$2\tau^{2}NM^{4}C^{2} +$$2\tau^{2}NM^{4}C^{2}$
Table 1: Multiplications in MSA on $\mathbf{X}\in\mathbb{R}^{N\times M^{2}\times C}$. $\tau<1$.
Computational cost comparison of the vanilla versus the proposed DCT-compressed attentions is presented in Table 1. We evaluate both naive and simplified implementations of the DCT-compressed attention with the vanilla attention. The simplification is implemented by fusing IDCT into the linear layer $\mathcal{L}_{O}$. Computational overhead from the Softmax function is omitted because the Softmax function is based on exponential terms. Despite this, the Softmax function in the DCT-compressed attention requires less computational cost than in the vanilla attention as the number of entries in $\left(\frac{\mathbf{QK}^{T}}{\sqrt{d}}+\mathbf{B}\right)$ in the DCT-compressed attention is $\tau^{2}$ times ($\tau<1$) as in the vanilla attention.
3.4 Frequency Domain Motivation of the DCT-based Initialization
In traditional random initialization, all the weight matrices are initialized as random numbers that are independently generated. As a result, the initial matrices can be considered as a realization of white noise covering the entire spectrum. A good initialization should take advantage of the full bandwidth to cover the entire spectrum of the input. The DCT-based initialization covers the entire frequency domain: The set of DCT basis vectors $\{\frac{\sqrt{2}}{2},\cos\frac{(2n+1)k\pi}{2C}\}$ is the class of discrete Chebyshev polynomials:
@@ 3.4 Frequency Domain Motivation of the DCT-based Initialization
for $n=0,1,\ldots,C-1$. The basis member $\cos\frac{(2n+1)k\pi}{2nC}$ is the $k$-th Chebyshev polynomial $T_{k}[\xi]$ evaluated at the $n$-th zero of $T_{n}[\xi]$ Ahmed et al. (1974). Fig. 2 draws the basis vectors in Eq. (6) for an $8\times 8$ DCT matrix and Fig. 3 presents the magnitude of their discrete Fourier transform ($\mathcal{F}\{\cdot\}$):
@@
Figure 3: Frequency response of the DCT basis vectors. Each frequency response is shown in a solid curve in different colors. The black dashed line denotes the root of the squared sum of these frequency responses, showing that the entire spectrum is covered.
3.5 Motivation of the DCT-based Compression
The mechanism behind the DCT-based compression mimics the Karhunen-Loeve transform (KLT), which is also known as principal component analysis (PCA), a mathematical technique used for dimensionality reduction in signal processing Ji et al. (2022). Neighboring image patches are highly correlated with each other. During training, the attention matrices are inter-correlated because they are obtained using similar gradient vectors. Therefore, they can be decorrelated and compressed using DCT. When the input data is highly correlated, the covariance matrix can be approximated by a Teoplitz matrix $\phi$ of $1,\rho,\rho^{2},\cdots,\rho^{C-1}]$:
@@ 3.5 Motivation of the DCT-based Compression
with $\rho<1$ close to 1 Akansu and Torun (2012). This is the covariance matrix of an autoregressive model random process AR(1) with the correlation coefficient $\rho$ Testa and Magli (2016). Fig. 2 shows that DCT basis vectors approximate the KLT of a first-order Markov process with high correlation among the neighboring samples Dony and others (2001); Radünz et al. (2022). We observe that $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ have parameters with high correlations among themselves. We can directly use DCT in data compression since its matrix approximates the corresponding KLT matrix Ahmed et al. (1974). Although KLT can also be utilized in data compression, it requires massive computation for the covariance matrices and the corresponding eigenvectors of $\mathbf{X}$. On the other hand, DCT is very successful in decorrelating the data without computing the covariance matrix and eigenvectors. As a consequence, it is also used in JPEG image Wallace (1991) and MPEG video Ebrahimi and Horne (2000) compression methods.
4 Experimental Results
@@
We conduct experiments on the CIFAR-10 and ImageNet-1K classification tasks and COCO 2017 detection and segmentation tasks. Swin Transformers Liu et al. (2021) were employed as the backbone network. Models were trained on a server with 8 NVIDIA A100 80GB GPUs using Python with the PyTorch framework.
4.1 Implementation Details
Image Classification
@@ 4.1 Implementation Details
Training models for the image classification tasks from scratch mostly followed the settings in Liu et al. (2021). We employed AdamW optimizer for 300 epochs using a cosine decay learning rate scheduler and 20 epochs of linear warm-up. We set the batch size as 1024, the initial learning rate as 0.001, and the weight decay as 0.05. Data augmentation was performed as follows: The CIFAR-10 augmentation followed He et al. (2016). First, the training images were padded with 4 pixels. Then, they were randomly cropped to get 32 by 32. Finally, the images were randomly flipped horizontally. The images were normalized with the means of [0.4914, 0.4822, 0.4465] and the standard deviations of [0.2023, 0.1994, 02010]. The ImageNet-1K augmentation followed PyTorch Torchvision’s GitHub repository Paszke et al. (2019). As a start, images were resized to make the smaller edge of the image 232 for Swin-T and 246 for Swin-S. Next, they were cropped to 224 by 224 images and randomly flipped horizontally. Techniques such as Autoaugment Cubuk et al. (2018), mixup Zhang et al. (2018), cutmix Yun et al. (2019), and random erasing Zhong et al. (2020) were applied. Images were normalized with the means of [0.485, 0.456, 0.406] and the standard deviations of [0.229, 0.224, 0.225]. During the training, the best models were stored based on the top-1 accuracy on the CIFAR-10 test dataset and the ImageNet-1K validation dataset.
COCO Detection and Segmentation
[…]
4.2 Image Classification Benchmark
Model Param(M) Top-1 Acc
Swin-T 27.53 90.51%
Swin-T-DCT-Q 27.53 90.78%
Swin-T-DCT-K 27.53 90.83%
Swin-T-DCT-V 27.53 90.81%
Swin-T-DCT-0.25 21.45 90.46%
Swin-T-DCT-0.5 22.67 90.67%
Swin-T-DCT-0.75 24.69 90.90%
@@ 4.2 Image Classification Benchmark
Table 2: CIFAR-10 test dataset evaluation. “-DCT-” followed by Q/K/V in model names indicates which attention projection weights are initialized with the DCT matrix, while numerical values denote the proportion of coefficients retained in DCT-compressed attention. Models with higher or lower top-1 accuracy are highlighted in pink or cyan, respectively.
CIFAR-10
@@
To demonstrate the benefits of using our DCT-based decorrelated attention on the small-scale dataset, we conduct Swin Transformer Tiny (Swin-T) from-scratch CIFAR-10 training. Table 2 presents the performance comparison on the CIFAR-10 test dataset, where the number of parameters only counts trainable parameters. The baseline vanilla Swin-T obtained 90.51% accuracy, and its accuracy was always improved by the DCT-based initialization regardless of which attention weight was initialized as the DCT matrix. When the weights to compute keys ($\mathbf{K}$) were initialized as the DCT matrices, the accuracy was improved to 90.83%. For DCT-compressed attention, when we only keep 25% of DCT coefficients, trainable parameters were reduced from 27.53M to 21.45M (22.1% off), while the accuracy only dropped 0.05%. If we kept more DCT coefficients, the accuracy even improved. 75% of DCT coefficients increased accuracy from 90.51% to 90.90%. This is because although the number of trainable parameters is reduced, transforming the input data into the frequency domain using DCT can boost the extraction of important features. As a result, fewer trainable parameters are required.
Method Param(M) Flops(G) Top-1 Acc Top-5 Acc
Swin-T 28.29 4.49 81.474% 95.776%
PoolFormer-S24 21 3.5 80.3% -
DCFormer-SW-T ($\tau=1)$ 28.29 4.5 81.2% -
DCFormer-SW-T ($\tau=2)$ 28.29 1.3 79.2% -
gSwin-T 22 3.6 81.715% -
@@
Swin-S 49.61 8.74 83.196% 96.360%
PoolFormer-M36 56 8.8 82.1% -
DCFormer-SW-S ($\tau=1)$ 49.61 8.7 82.8% -
DCFormer-SW-S ($\tau=2)$ 49.61 2.7 80.9% -
gSwin-S 40 7.0 83.014% -